### v1

# The Illusive Slump of Disruptive Patents

Abstract
Despite tremendous growth in the volume of new scientific and technological knowledge, the popular press has recently raised concerns that disruptive innovative activity is slowing. These dire prognoses were mainly driven by Park et al. (2023), a Nature publication that uses decades of data and millions of observations coupled with a novel quantitative metric (the CD index) that characterizes innovation in science and technology as either consolidating or disruptive. We challenge the Park et al. (2023) methodology and findings, principally around concerns of truncation bias and exclusion bias. We show that 88 percent of the decrease in disruptive patents over 1980-2010 reported by the authors can be explained by their truncation of all backward citations before 1976. We also show that this truncation bias varies by technology class. We update the analysis to 2016 and account for a change in U.S. patent law that allows for citations to patent applications in addition to patent grants, which is ignored by the authors in their analysis. We show that the number of highly disruptive patents has increased since 1980—particularly in IT technologies. Our results suggest caution in using the Park et al. (2023) methodology as a basis for research and decision making in public policy, industry restructuring or firm reorganization aimed at altering the current innovation landscape.
Keywords: Disruptive Innovation, Truncation Bias, Exclusion Bias, U.S. Patent Law Change
JEL: 030, 032, O33
1 Introduction
Continued innovation in science and technology is considered a bedrock for and driving force of growth and prosperity in most economies. A paper co-authored by Park, Leahey and Funk entitled, “Papers and Patents are Becoming Less Disruptive Over Time” was recently published in Nature (2023). Given its provocative title and findings, as well as the topic examined, the paper attracted significant and global media attention. Park et al. (2023) has been featured in hundreds of international newspapers and magazines Altmetric (2023): for example, The Economist (2023) emphasized in a report on the changing nature of science that "Papers and Patents are Becoming Less Disruptive", indicating in a subtitle that "why that is, is a mystery"; The New York Times (2023) further stated: "What Happened to All of Science’s Big Breakthroughs?"; and the Financial Times (2023) noted: "Science is Losing its Ability to Disrupt".
One reason for this attention is that the Park et al. (2023) results show marked declines in disruptive innovation over time. The authors state in the abstract that their "results suggest that slowing rates of disruption may reflect a fundamental shift in the nature of science and technology" (p. 138), and note in the conclusion that "this trend is unlikely to be driven by changes in citation practices or the quality of published work. Rather, the decline represents a substantive shift in science and technology, one that reinforces concerns about slowing innovative activity. We attribute this trend in part to scientists’ and inventors’ reliance on a narrower set of existing knowledge" (p. 142).
Another reason for this attention is that the Park et al. (2023) results have implications regarding the organization of the entire science and technology innovation process—from government research labs and universities to private and public enterprises. The authors’ methodology and findings unsurprisingly piqued the interests of researchers, commentators, and journalists on a global scale who offered myriad explanations, such as increased pressures on scientists to apply for large (interdisciplinary) projects, rising administrative burdens, declining basic research funding, increasing risk-aversion among scientists, and pressures to publish rapidly Rust (2023); Yanai and Lercher (2023). Still other researchers raised concerns around the Park et al. (2023) methodology as a measure of disruptive innovation Ruan et al. (2021); Bornmann et al. (2020a, b).
An important question arises, however, as to whether the argument made by Park et al. (2023) is accurate. We have major concerns around their methodological approach regarding truncation and exclusion bias related to measurement. We show using patent data that the worrisome result suggested by the authors is instead mainly a consequence of the omission of citations to older innovations: i.e., the authors artificially truncate all backward citations of patents before 1976. We also show using patent data that substantive differences in the results occur from not considering patent law changes: i.e., the authors neglect backward citations of patents published after 2000 to patent applications. Both of these biases have large measurement effects on their proposed CD index and direct implications on the accuracy of their findings and conclusions.
In what follows, we first compare the Park et al. (2023) methodology with truncation to our own methodology with no truncation. We explain why truncation bias can have such a significant impact, and then assess the effects of truncation bias using patent data. We then update the patent data over 2011-2016 and take into account a change in patent law that allows for citations to patent applications as well as to patent grants. We show that the number of highly disruptive patents has increased since 2005. We conclude with a discussion of our main findings and the dangers of measurement bias in influencing public policy, industry organization, and firm organization decisions and interventions.
2 Measuring Patent Disruptiveness
The CD index was developed by Funk and Owen-Smith (2017) and used by Park et al. (2023) to characterize whether a patent or scientific publication is considered more consolidating (i.e., building upon previous research and reinforcing the status quo) or more disruptive (i.e., obsolescing previous research and pushing into new directions). [^1] Since its introduction by Funk and Owen-Smith (2017) the index has been used in a wide variety of analyses to capture the disruptive content of patents (e.g. Kaltenberg et al., 2023; Wu et al., 2019) and of scientific papers (e.g. Bornmann et al., 2020a; Wu et al., 2019). It is based on the idea that a patent or scientific publication is disruptive if "the subsequent work that cites it is less likely to also cite its predecessors" (Park et al., 2023, p. 139). The CD index ranges from -1 (consolidating) to 1 (disruptive). The measure uses five-year post-publication windows in its construction, referred to as $CD_{5}$. For scientific papers, it starts in 1945; for patents, it starts in 1980.
Fig. 1 plots two average $CD_{5}$ indices for approximately $3.66$ million patents over $1980$-$2010$ from the United States Patent and Trademark Office (USPTO): one based on the Park et al. (2023) methodology; the other based on our methodology. [^2] As in Park et al. (2023), we consider only utility patents and draw our data from PatentsView (version February $21$, $2023$). The database contains all US patents that have been published between $1976$ and September $29$, $2022$. While Park et al. (2023) focus on patents of the aggregate NBER technology fields ”Chemical”, ”Computer and Communications”, ”Pharmaceutical and Medical”, ”Electrical and Electronic”, and ”Mechanical”, which encompasses a total of $3,046,672$ granted utility patents over 1980-2010, our main analysis utilizes all granted utility patents published by the USPTO during the same period, resulting in a total of $3,662,051$ patents. For more details, see Appendix A. As is readily apparent, the evolution of each index differs markedly: the Park et al. (2023) methodology produces an average annual CD index that starts at 0.39 in 1980 (the first observation year) but declines rapidly over time; our methodology instead indicates an average annual CD index that starts at 0.09 in 1980 and declines more gradually over time. Fig. 1 nonetheless shows marked convergence in the two indices in the later years of the sample—particularly since 2000.
Figure 1: The CD index: Park et al. (2023) Methodology Versus Our Methodology
The figure shows the average $CD_{5}$ Index calculated using the Park et al. (2023) methodology and our methodology. The Park et al. (2023) methodology (in blue) omits backward citations to granted USPTO patents published before 1976; our methodology (in yellow) includes all backward citations to granted USPTO patents.
The difference in the average CD indices arises from Park et al. (2023) truncating all backward citations before 1976. To the best of our understanding, the authors do not explain or acknowledge this truncation explicitly despite it having important consequences. The Park et al. (2023) truncated CD index has much larger values compared to our untruncated CD index for those patents published closer to the truncation year. The reason is simple: patents published closer to the truncation year (e.g., 1980) frequently cite patents published in proximate years (e.g., pre-1976), while patents published well-past the truncation (e.g., 2010) infrequently cite those patents. In the earlier years, many backward citations are truncated; in the later years, only a few are. As a result, the CD indices diverge in the early years and converge in the latter years.
This implies that the Park et al. (2023) results are significantly biased due to the truncation of backward patent citations, with the bias strongest for those patents published closer to the truncation year. Hence, their findings of a sharp decline in the average patent disruptiveness can mostly be attributed to a reduction in measurement bias; not to a decline in disruptive innovations. After correcting for truncation bias, our methodology shows that the CD index has a much lower starting point and more modest decline over time. Specifically, 88 percent of the total decline in the average disruptiveness of patents found by Park et al. (2023) can be explained by the authors artificially truncating all backward citations to patents published before 1976.
3 Why Truncation Matters
An example helps illustrate our argument: We compare two patents that both appear in the Park et al. (2023) dataset: [^3] Fig. B.3 in the Appendix shows the first page from each patent document listing the backward citations. US 4181011 was issued in 1980; US 6511791 was issued in 2003. US 4181011 makes 13 citations to patents published between 1958 and 1974; US 6511791 makes 11 citations to patents published between 1987 and 2001. Using the Park et al. (2023) methodology and replication data ($patentsview\_analytical\_df.csv$), the number of backward citations for patent US 4181011 is zero—given the exclusion of backward citations to patents published before 1976. In contrast, the number of backward citations for patent US 6511791 is 11—the same number as listed in the patent document.
Figure 2: How Truncation of Backward Citations Affects the CD Index
The illustration shows how the CD index of a patent is biased upwards if backward citations are truncated.
Fig. 2 provides intuition as to how the truncation of backward citations at a given point in time can bias the CD index: without truncation, the focal patent has a CD index value of -0.5 and is considered consolidating; with truncation (i.e., excluding pre-1976 backward citations), the focal patent has a CD index value of 1 and is considered disruptive.
Formally, this can be shown as follows. Park et al. (2023) define the CD index for a focal patent as:
$CD_{t}=\frac{1}{N}\sum_{N=i}^{N}(-2f_{it}b_{it}+f_{it}),\leavevmode\nobreak\ with$ (1)
$f_{it}=\begin{cases}1,&\text{if $i$ cites the focal patent within $t$ years ofpost-publication of the focal patent}\\ 0,&\text{otherwise}\end{cases}$
and
$b_{it}=\begin{cases}1,&\parbox[t]{346.89731pt}{if $i$ cites parts of the predecessors of the focal patent within $t$ years of post-publication of the focal patent}\\ 0,&\text{otherwise}\end{cases}$
where $N$ is the sum of the number of patents citing: the focal patent only ($N_{F}$), the focal patent and parts of its predecessors ($N_{B})$, and parts of its predecessors only ($N_{R}$) within $t$ years after the focal patent publication. A forward citation increases the consolidating (disruptive) nature of the focal patent if the forward cited patent does (does not) cite the predecessor patents (i.e., those patents also cited by the focal patent).
We extend the CD index by directly considering truncation. Truncation affects the CD index in two ways. First, left-truncation (i.e., with respect to the earlier years) reduces the number of counted forward citations $N$ if the truncated backward citations are cited within the $t$ post-publication years of the focal patent:
$N^{tr}<N^{non-tr},$ (2)
where $tr$ stands for truncated and $non-tr$ for non-truncated backward citations.
Second, left-truncation reduces the $b_{it}$ value if a forward citation of the truncated predecessor is also a forward citation of the focal patent within $t$ years of publication. In this case:
$-f_{it}b_{it}^{tr}=0\text{ and }-f_{it}b_{it}^{non-tr}=-1\rightarrow-f_{it}b_{it}^{tr}>-f_{it}b_{it}^{non-tr}.$ (3)
Considering equations (2) and (3), we distinguish among four distinct cases: (i) no backward citations are truncated; (ii) backward citations are truncated but none are cited within the $t$ post-publication years of the focal patent; (iii) truncated backward citations are cited within the $t$ post-publication years of the focal patent but none cite the focal patent; and (iv) the truncated backward citations are cited within the $t$ post-publication years of the focal patent and at least one of them cite the focal patent. In case (i) and case (ii), equations (2) and (3) hold as equalities and truncation has no effect on the CD index:
$CD_{t}^{tr}=CD_{t}^{non-tr}.$ (4)
In case (iii), the inequality of equation (2) holds but equation (3) is strictly an equality. In case (iv), the inequalities of equations (2) and (3) hold. In case (iii) and case (iv), the truncated CD index is biased upward as long as the summation part of equation (1) of the truncated CD index is positive, because each of the two effects artificially increases its value. Moreover, ceteris partibus, the bias is stronger in case (iv). Both results are easily seen by adding the inequalities of equations (2) and (3) to the CD index provided in equation (1). As a result:
$CD_{t}^{tr}>CD_{t}^{non-tr}.$ (5)
Finally, truncation may lead to an upward or downward biased CD index if the summation of equation (1) is negative and the inequalities of either equation (2) or equations (2) and (3) hold. In this case:
$CD_{t}^{tr}\gtrless CD_{t}^{non-tr}.$ (6)
4 Assessing Truncation Effects
Fig. 3 categorizes focal patents by the number of backward citations to demonstrate the impact of truncation. The left panel uses the Park et al. (2023) methodology; the right panel uses our methodology.
Figure 3: Patent Percentage by Backward Citation Categories
Notes: The figure divides patents into five categories based on the number of backward citations. The vertical-axis shows the percentage of focal patents in each category. The left panel (Park et al. (2023) Methodology) excludes backward citations to patent grants published before 1976; the right panel (No Truncation Methodology) includes backward citations to patent grants regardless of publication year.
Both panels show that the majority of patents have backward citations. However, notable differences are found depending upon whether truncation is or is not present. The left (truncation) panel indicates that 38 percent of patents had no backward citations in 1980 and 99 percent of patents had no backward citations in 1976 (the truncation year), given that the only backward citations included are those patents also published in 1976. The right (no truncation) panel indicates that less than two percent of patents had no backward citations in these years. Fig. 3 also indicates that patents published further from the truncation year have lower proportions of truncated backward citations. The reason is again straightforward: More recent patents tend to cite more recent patents.
Figure 4: Truncation Time is Not Important
The figure shows average $CD_{5}$ indices excluding backward citations to patents published before 1976 (blue line) as in Park et al. (2023), before 1986 (green line), and before 1996 (black line). Park et al. (2023) use a four-year window from the truncation year for the start of their analysis. The dashed lines show how the truncated $CD_{5}$ indices change post-truncation year and within this four-year window for illustrative purposes. The solid lines show how the $CD_{5}$ indices change post-truncation year and beyond this four-year window (as examined by Park et al., 2023). The average $CD_{5}$ index with no truncation is included for comparison purposes.
It is readily apparent that the number of years considered for backward citations matters. The analysis of Park et al. (2023) begins in 1980 and includes backward citations to 1976 (the first year that USPTO patents are available). Hence, patents published in 1980 include four years of backward citations. Fig. 4 shows that had their analysis started in 1976 (i.e., the year of truncation), the bias would be much larger: nearly all patents from 1976 would be classified as disruptive. Fig. 4 also shows that the starting year decision does not alleviate truncation bias and measurement concerns: e.g., had their analysis started in 1990 or in 2000, a four-year window for backward citations would produce similarly sharp declines in the CD index in the earliest years after the truncation year which then stabilizes and converges to the unbiased index over time.
Fig. 5 indicates that this bias is not identical across the technologies examined by Park et al. (2023). The bias is most pronounced in the aggregate technology "Mechanical" and least pronounced in "IT". The difference is again most likely explained by truncation: IT patents have developed more recently and rapidly in comparison to Mechanical patents, suggesting older patents are relatively more often cited in the latter category than in the former category (see Appendix Fig. B.4). As a result, fewer backward citations are truncated in IT patents and more backward citations are truncated in Mechanical patents.
Figure 5: Size of Truncation Bias Differs by Technologies
The figure shows the average $CD_{5}$ index for various aggregate WIPO technologies, as determined in Table B.3. The blue line shows the CD index when backward citations to granted U.S. patents published before 1976 are excluded, per the Park et al. (2023) methodology. The yellow line shows the $CD_{5}$ index when all backward citations are included, per our methodology.
5 Updating Years and Patent Law Change
We examine two factors that could affect the CD index post-2010. First, we update the analysis window using patent data to 2016. Second, we consider a major change to U.S. patent law: in particular, the Inventor Protection Act of 1999 requires that U.S. patent applications be published 18 months after the initial application filing, effectively shifting citation patterns from strictly patent grants to both patent grants and patent applications Johnson and Popp (2003). [^4] See Appendix A for more details on the additional data used, and Appendix B for a more detailed discussion of how excluding citations to patent applications affects the CD index of patents. Fig. 6 shows these post-2010 changes in three average $CD_{5}$ indices: the Park et al. (2023) methodology (I); our methodology including all backward citations to strictly patent grants (II); and our methodology including all backward citations to patent grants and patent applications (III). Methodology (III) considers only those citations to patent applications that receive patent grants by the end of 2021. This approach best achieves consistency between the pre- and post-Inventor Protection Act periods, as only patent grants were allowable citations. [^5] Fig. B.5 in the Appendix shows similar results when citations to patent applications that have not (yet) received patent grants are included.
Figure 6: Patent Law Change Creates Exclusion Bias in Park et al. (2023) Methodology
The figure shows the average $CD_{5}$ index under different methodologies: the Park et al. (2023) methodology that excludes backward citations to patent grants before 1976 (I); our methodology that includes all backward citations to strictly patent grants (II); and our methodology that includes all backward citations to patent grants and patent applications (III).
Fig. 6 is identical to Fig. 1 up to 2001 because—prior to the Inventor Protection Act’s passage—patents only cited patent grants. The inclusion of citations before 1976 eliminates truncation bias and subsequently decreases the CD index over 1980-2005—as seen by comparing (I) to (II) and (III). The inclusion of citations to patent applications eliminates exclusion bias and subsequently decreases the CD index over 2005-2016—as seen by comparing (I) and (II) to (III). Correcting for both biases indicates that the average CD index decreases slightly over 1980-2016 and is relatively stable since 2005—as shown in methodology (III).
6 Increasing Number of Highly Disruptive Patents
[…]
Figure 7: Investigation of Highly Disruptive Patents
(a) Different Calculation Methods
(b) Method (III) and Different Technologies
The figure illustrates the number of and change in highly disruptive patents (i.e., $CD_{5}\in(0.75,1]$) over time. Panel (a) shows the number of highly disruptive patents under different methodologies: the Park et al. (2023) methodology that excludes backward citations to patent grants before 1976 (I); our methodology that includes all backward citations to strictly patent grants (II); and our methodology that includes all backward citations to patent grants and patent applications (III). The total number of highly disruptive patents differs from that reported in Fig. 4b (p. 141) of Park et al. (2023) because they only consider a subset of technologies. (See Appendix A for more detailed discussion.) Panel (b) shows the evolution of highly-disruptive patents across aggregated WIPO technology categories using methodology (III), normalized using 1980 as a baseline.
Using 1980 as a benchmark, Panel (b) of Fig. 7(b) shows that disruption is heterogeneous across technologies: the number of highly-disruptive IT patents increases nearly twenty-fold; the number of highly-disruptive Pharma patents increases more than four-fold; and the number of highly-disruptive Mechanical patents is stable. These results challenge the Park et al. (2023) argument that "despite large increases in scientific productivity, the number of papers and patents with $CD_{5}$ values in the far right tail of the distribution remains nearly constant over time" (p. 140). Our findings are, however, consistent with a recent scientific publication that indicates "a paper published in Nature Park et al. (2023) showed that research is becoming less disruptive. This does not seem to be the case for the next generation of cancer chemotherapy based on the methionine dependence of cancer" (p. 277, Kubota et al., 2023).
Thus far, we have taken an aggregated methodological approach, but this masks two disparities inherent at the individual patent level. First, the truncation of backward citations misclassifies a large proportion of patents as highly disruptive. Appendix Table B.1 illustrates how $CD_{5}$ categories for patents published in 1980 calculated using the Park et al. (2023) (truncation) methodology are distributed across $CD_{5}$ categories using our (no truncation) methodology. As an example, 25 percent of the USPTO patents in 1980 that are considered highly disruptive under the truncation methodology are actually consolidating under the no truncation methodology.
Second, the exclusion of backward citations to patent applications misclassifies a large proportion of recent patents as highly disruptive. Appendix Table B.2 shows how patents in 2016 that belong to a particular $CD_{5}$ category using the Park et al. (2023) (truncation) methodology are distributed across $CD_{5}$ categories using our (no truncation) methodology and accounting for backward citations to patent applications. As an example, 19 percent of the USPTO patents in 2016 that are considered highly disruptive using the Park et al. (2023) methodology are actually consolidating when accounting for both truncation and backward citations to patent applications. Determining the disruptive potential of individual patents thus requires careful consideration to ensure against bias and misclassification. Such bias and misclassification is the case when the data provided by Park et al. (2023) or Funk and Owen-Smith (2017) are used for analysis and decision-making.
7 Discussion
Our examination of the Park et al. (2023) CD index methodology was prompted by the stark decrease in the share of disruptive innovation in science and technology the authors find in patents over 1980-2010 and in scientific publications over 1945-2010. At least for patents, our analysis instead suggests that the decrease in the CD index is mainly due to the truncation of backward citations, while a modest increase in the CD index post-2005 is mainly driven by the exclusion of backward citations to patent applications brought on by a change in patent law.
Variations of the index proposed in Park et al. (2023) and based upon other research Leydesdorff et al. (2021); Bornmann et al. (2020a) are also affected by truncation bias and exclusion bias when backward citations to granted patents and to patent applications are not appropriately accounted for. Our analysis is important for the research community that may overlook these fundamental issues and work with biased data – either via the CD index or some other measure.
Our analysis highlights, in particular, the dangers of truncation bias from backward citations and exclusion bias from citations to patent applications. Our correction and methodology shows—in comparison to Park et al. (2023)—that: (1) the CD index starting level is significantly lower, modestly declines up to 2005, and is mostly stable since; (2) the starting truncation "date" does not matter—any backward citation "cutoff" in any chosen year can produce similar CD index patterns; (3) truncation bias differences are heterogeneous across technology classes; and (4) the number of highly disruptive patents has increased—particularly since 2010 and in IT.
Our analysis suggests, in general, caution against overt government policy interventions, research policy changes, or substantive industry and firm reorganizations that seek to improve or disrupt the innovation status-quo. If it was accurate that breakthroughs in science and technology have substantially decreased over time, the current organization of the entire science and technology innovation process could be called into question and should be reevaluated. In particular, reforms that seek to improve the creation of disruptive innovation or restart innovation within R&D centers and research labs—e.g., in government agencies, universities and firms—would become paramount.
But our results suggest—at least in the patent data—that the so-called recent slump in disruptive innovation is illusory. An unbiased CD index is instead markedly stable over the past several decades, and highly disruptive patents have increased overall–especially in IT and Pharma. Thus, an interesting question for future research could investigate why these industries have seen increases in disruptive innovation, relative to others. Given our findings, we should finally emphasize that premature changes in the current innovation processes of science and technology might bring with them unintended consequences.
8 Credit Authorship Contribution Statement
Jeffrey T. Macher: Conceptualization, Writing – original draft, Writing – review & editing. Christian Rutzer: Calculation, Conceptualization, Investigation, Visualization, Writing – original draft, Writing – review & editing. Rolf Weder: Conceptualization, Writing – original draft, Writing – review & editing.
9 Declaration of Competing Interest
The authors have received no external funding and declare no competing interests.
10 Data Availability
The code and data needed to reproduce this study are available at Zenodo https://zenodo.org/deposit/8020004. The raw data used in this research was obtained from PatentsView, a platform that provides unrestricted access to detailed information on all granted USPTO patents published since 1976. All raw data can be downloaded at https://patentsview.org/download/data-download-tables.
We used R version 4.1.2 to perform the computations, mainly using the tidyverse and data.table libraries for data processing and ggplot2 for visualization. The computations were performed using SLURM on the sciCORE cluster at the University of Basel, Switzerland.
A Data
The study uses data from the February 21, 2023 version of PatentsView, which includes all U.S. patents published between 1976 and September 29, 2022. As in previous research Park et al. (2023); Funk and Owen-Smith (2017), we focus only on USPTO patents. We include PatentsView data that assigns each USPTO patent to technology categories as defined by the World Intellectual Property Organization (WIPO) Schmoch (2008). The specific WIPO technologies used to create our aggregated technology groups are provided in the Table B.3. It is important to note that our technology classification differs from that of Park et al. (2023), who categorize patents based on NBER technology fields that are no longer available.
In Sections 2-4, we use all granted utility patents over 1980-2010. Park et al. (2023) examine granted utility patents over this timeframe and in the NBER technology fields of "Chemical," "Computers and Communications," "Drugs and Medical," "Electrical and Electronic," and "Mechanical." Their data includes 3,046,672 patents and 29,777,375 backward citations (see replication data file $patentsview\_analytical\_df.csv$ of Park et al. (2023)). We examine granted utility patents over the same timeframe, but cover all technology fields. Our data thus includes 3,662,051 patents and 42,617,016 backward citations in the case of no truncation (i.e., as they appear in the original patent documents) and 35,192,186 backward citations in the case of truncation.
In Sections 5-6, we expand the dataset over 1980-2016, which results in 5,324,224 granted utility patents and 74,943,850 backward citations. With the passage of the Inventor Protection Act of 1999, patents beginning in November 2000 include backward citations to patent applications and to patent grants. Thus, in our methodology (III), we include all backward citations to patent grants and to patent applications that were eventually granted using supplemental PatentsView data. We replace all citations to patent applications with the corresponding patent grant information: i.e., the grant number and publication date. This replacement can result in some backward citations with publication dates later than the citing patent, but the backward citation publication date is not relevant to the $CD_{5}$ index methodology. What matters are publication dates of forward citations. We do not include patent applications published within five years after the publication year of a focal patent but with corresponding patent grants published later in the analysis. Finally, we do not include citations to patent applications that have not been granted by the end of 2021 (the latest date available) in the analysis. This approach results in 87,880,452 backward citations.
B CD Index and Patent Law Change
The Inventor Protection Act of 1999 might bias the CD index if citations to patent applications are not considered—especially since patent applications are now increasingly cited vis-a-vis patent grants. As Kuhn et al. (2020) note, citations to patent applications accounted for 25 percent of all citations made by USPTO patents in 2015.
Figure B.1: Excluding Backward Citations to Patent Applications Affects the CD Index
The illustration shows how excluding citations to patent applications can bias the CD index of a patent, similar to the truncation of backward citations shown in Figure 2.
First, as Fig. B.1 illustrates, ignoring patent applications can create bias by falsely declaring a patent as completely disruptive. The channels are the same as when backwards citations are truncated in time. Formally, it affects the CD index via $b_{it}$ and $N$ in equation (1), but the driving force is now missing citations to patent applications. As shown by equations (2)-(6), this can create a bias in the CD index similar to left-truncated backward citations.
Figure B.2: Backward Citations to Patent Application and Corresponding Patent Grant Affects the CD Index
The illustration shows backward citations made to a patent application and the corresponding patent grant can bias the CD index. The patent can be mis-classified as fully disruptive due to the exclusion of citations to the patent application. Specifically, the analysis misses citation to a predecessor application document as well as citation to the application document of the focal patent.
Second, ignoring patent applications can create bias if a patent application is cited in one instance and the subsequent patent grant is cited in another instance. This can occur from citations to predecessor patents or to the focal patent—i.e., applications and grants. Fig. B.2 illustrates this possibility in more detail. Formally, it affects the CD index in equation (1) as before in $b_{it}$ or $N$ but now also in $f_{it}$. This may result in an upward or downward bias of the CD index. An upward bias occurs, for example, when the focal patent cites a patent application and a successor patent cites the focal patent and the patent grant (instead of the patent application). In such a case, the citation to the patent application is missed. A downward bias occurs, for example, if a successor patent only cites the patent application of the focal patent and no predecessor patents (applications or grants) of the focal patent. As this would not be considered a forward citation to the focal patent, it thereby loses some of the disruptive value of the focal patent. The Inventor Protection Act of 1999 may therefore bias the CD index if citations to patent applications for patents published after November 29, 2000, are not properly taken into account.
B.1 Figures
Figure B.3: Examples of Backward Citations in Patent Documents
(a) First Page of US Patent 4181011
(b) First Page of US Patent 6511791
The left panel shows US Patent 4181011 makes 13 backward citations to granted patents, all of which were published before 1976. The right panel shows US Patent 6511791 makes 11 backward citations to other granted patents, all of which were published after 1976.
Figure B.4: Average Backward Citation Age
The figure shows the average age of backward citations for various aggregate WIPO technologies, as determined in Table B.3. The left panel (Truncation Methodology) shows that the average backward citation ages increase in each technology because citations before 1976 are not considered. The middle panel (No Truncation Methodology) shows average backward citation ages are relatively stable in each technology up to 2005 and then increase markedly post-2005 because citations to patent applications are not considered. The right panel (No Truncation + Patent Application Citation Methodology) shows average backward citation ages are relatively stable in each technology and over time.
Figure B.5: Average CD Index By Various Methods
The figure shows the average CD index calculated using four different methodologies. Methodology (I) via Park et al. (2023) excludes backward citations to patent grants before 1976 and backward citations to patent applications. Methodology (II) includes backward citations to patent grants before 1976, but excludes backward citations to patent applications. Methodology (III) includes backward citations to patent grants before 1976 and to patent applications that have received patent grants by the end of 2021. This methodology replaces the application number with the corresponding patent grant number and publication date of the application with the publication date of the patent grant. Finally, methodology (IV) includes backward citations to patent grants before 1976, to patent applications that have received patent grants by the end of 2021, and to patent applications that have not received patent grants by the end of 2021. This methodology replaces the application numbers with their corresponding granted patent numbers while keeping the original application dates, which is necessary to ensure that all citations are taken into account as illustrated by Fig. B.2. Note methodology (IV) lies slightly below methodology (III) after 2005 due to the inclusion of backward citations to patent applications that did not receive patent grants and thus would not appear prior to the Act’s passage. Citations to patent applications can increase the number of citations to predecessor patents, which in turn may lead to a decrease in the CD index as the value of $N$ in equation (1) is larger. In this case, however, the decline in the CD index does not necessarily suggest a decline in the number of disruptive innovations. Instead, it could be due to what Petersen et al. (2019) terms "citation inflation".
B.2 Tables
Table B.1: Truncated vs. Non-Truncated Distribution of Patents’ $CD_{5}$ Index for 1980
Truncation
No Truncation -1, 0 0, 0.25 0.25, 0.5 0.5, 0.75 0.75, 1
-1, 0 1.00 0.18 0.19 0.19 0.25
0, 0.25 0.00 0.82 0.52 0.34 0.33
0.25, 0.5 0.00 0.00 0.29 0.32 0.22
0.5, 0.75 0.00 0.00 0.00 0.16 0.05
0.75, 1 0.00 0.00 0.00 0.00 0.15
  • 1
    The table shows how the $CD_{5}$ values of patents of a particular group calculated using the Park et al. (2023) methodology are distributed across the $CD_{5}$ values using the correct methodology without truncation. Reading example: The last row of the last column shows that only 15 percent of all patents with a $CD_{5}$ index between 0.75 and 1 and calculated by the Park et al. (2023) methodology actually belong to the same group when all backward citations are taken into account.
Table B.2: Truncated vs. Non-Truncated Distribution of Patents’ $CD_{5}$ Index for 2016
Truncation
No Truncation and Patent application -1, 0 0, 0.25 0.25, 0.5 0.5, 0.75 0.75, 1
-1, 0 0.69 0.26 0.18 0.18 0.19
0, 0.25 0.29 0.71 0.63 0.52 0.56
0.25, 0.5 0.01 0.02 0.15 0.18 0.10
0.5, 0.75 0.00 0.00 0.03 0.10 0.04
0.75, 1 0.00 0.00 0.01 0.02 0.12
  • 1
    The table shows how patents belonging to a particular group of $CD_{5}$ values calculated with the Park et al. (2023) methodology are distributed across the $CD_{5}$ values using the correct methodology without truncation and considering citations to patent applications. Reading example: The last row of the last column shows that only 12 percent of all patents with a $CD_{5}$ index between 0.75 and 1 calculated by the Park et al. (2023) methodology belong to the same group when all backward citations and citations to patent applications are taken into account.
Table B.3: Aggregation of Technologies
WIPO Field ID WIPO Field Name Aggregate Technology
1 Electrical machinery, apparatus, energy Electronics
2 Audio-visual technology Electronics
3 Telecommunications Electronics
4 Digital communication IT
5 Basic communication processes Electronics
6 Computer technology IT
7 IT methods for management IT
8 Semiconductors IT
9 Optics Instruments
10 Measurement Instruments
11 Analysis of biological materials Chemical
12 Control Instruments
13 Medical technology Pharma
14 Organic fine chemistry Chemical
15 Biotechnology Chemical
16 Pharmaceuticals Pharma
17 Macromolecular chemistry, polymers Chemical
18 Food chemistry Chemical
19 Basic materials chemistry Chemical
20 Materials, metallurgy Chemical
21 Surface technology, coating Chemical
22 Micro-structural and nano-technology Chemical
23 Chemical engineering Chemical
24 Environmental technology Chemical
25 Handling Mechanical
26 Machine tools Mechanical
27 Engines, pumps, turbines Mechanical
28 Textile and paper machines Mechanical
29 Other special machines Mechanical
30 Thermal processes and apparatus Mechanical
31 Mechanical elements Mechanical
32 Transport Mechanical
33 Furniture, games other
34 Other consumer goods other
35 Civil engineering other
  • 1
    The WIPO technologies are defined in Schmoch (2008). The USPTO assigns each patent to one or more of these technologies. The assignment can be found in the data available in PatentsView.